# paper=a4paper,margin=2.0cm

### v1

@@ 1 Introduction
In addition to the above, rogue waves, a large localized wave structure that appears and vanishes without a trace [45], and breathers, which are localized breathing structures, have also been investigated in the LSRI system both in the $(1+1)$- and $(2+1)$-dimensional settings, from single-component to multicomponent systems, revealing interesting dynamical patterns [31, 32, 33, 34, 35, 36, 37, 38]. It should be noted that the variants of the LSRI model, such as the Newell (N) type LSRI system [39] and the generalized or N-YO type LSRI system [40], have been proposed. The various nonlinear wave solutions for these models have been reported in the literature [41, 42, 43, 44]. Considering all these developments, in this paper, we intend to investigate the recently proposed non-degenerate vector solitons [46, 47, 48, 49, 50, 51], characterized by non-identical wave numbers, and their several novel features in the following two-component LSRI system
@@ 2 Non-degenerate fundamental soliton solution and its special features
$e^{R_{1}}=-\frac{i\sigma_{1}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}+k_{1}^{*})^{2}(k_{1}-k_{1}^{*})},~{}e^{R_{2}}=-\frac{i\sigma_{2}|\alpha_{1}^{(2)}|^{2}}{2(l_{1}+l_{1}^{*})^{2}(l_{1}-l_{1}^{*})},$
$e^{\Delta_{1}^{(1)}}=-\frac{i\sigma_{2}\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|^{2}(k_{1}-l_{1})}{2(k_{1}+l_{1}^{*})(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}},~{}e^{\Delta_{1}^{(2)}}=\frac{i\sigma_{1}\alpha_{1}^{(2)}|\alpha_{1}^{(1)}|^{2}(k_{1}-l_{1})}{2(k_{1}^{*}+l_{1})(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}},$
$e^{R_{3}}=-\frac{\sigma_{1}\sigma_{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|k_{1}-l_{1}|^{2}}{4|k_{1}+l_{1}^{*}|^{2}(k_{1}-k_{1}^{*})(l_{1}-l_{1}^{*})(k_{1}+k_{1}^{*})^{2}(l_{1}+l_{1}^{*})^{2}}=e^{R_{1}+R_{2}+\Lambda},~{}e^{\Lambda}=\frac{|k_{1}-l_{1}|^{2}}{|k_{1}+l_{1}^{*}|^{2}}.$ (2d)
@@
$S^{(l)}(x,t)=\frac{\alpha_{1}^{(l)}e^{\eta_{1}}}{1+e^{\eta_{1}+\eta_{1}^{*}+R}}=2A_{l}k_{1R}\sqrt{k_{1I}}e^{i(\eta_{1I}+\frac{\pi}{2})}\mbox{sech}(\eta_{1R}+\frac{R}{2}),~{}A_{l}=\alpha_{1}^{(l)}(\sum_{l=1}^{2}\sigma_{l}|\alpha_{1}^{(l)}|^{2})^{-\frac{1}{2}},$ (3a)
$L(x,t)=2\frac{\partial^{2}}{\partial x^{2}}\ln(1+e^{\eta_{1}+\eta_{1}^{*}+R})=2k_{1R}^{2}\mbox{ sech}^{2}(\eta_{1R}+\frac{R}{2}),~{}R=\ln\frac{-\sum_{l=1}^{2}\sigma_{l}|\alpha_{1}^{(l)}|^{2}}{16k_{1R}^{2}k_{1I}},~{}l=1,2.$ (3b)
The above conventional vector bright soliton is described by only three arbitrary complex parameters, $\alpha_{1}^{(l)}$, $l=1,2$, and $k_{1}$. We remark that one can also obtain the solution (2a)-(2d) from the solution of degenerate two-solitons of the kind (3a)-(3b) in a non-standard way. The degenerate two-soliton solution can be deduced from Eqs. (25a)-(25d)) by setting $N=2$ and by restricting $l_{1}=k_{1}$ ($\xi_{1}=\eta_{1}$) and $l_{2}=k_{2}$ ($\xi_{2}=\eta_{2}$). By substituting $\alpha_{2}^{(1)}=\alpha_{1}^{(2)}=0$ in such degenerate two-soliton solution and renaming the constants, $\alpha_{2}^{(2)}$ and $k_{2}$ as $\alpha_{1}^{(2)}$ and as $l_{1}$, respectively, the solution (2a)-(2d) can be deduced. We have discussed this point in our earlier works in the case of Manakov system [47] and in the case of LSRI system (1) with $\sigma_{1}=\sigma_{2}=1$ [20] . However, in principle, it is possible to derive the non-degenerate soliton solution (2a)-(2d) systematically, see Appendix A of the present work and the methodology described in our earlier work [46, 47, 48, 49, 51, 50]. The dynamical features of such a new class of vector soliton solution (2a)-(2d) obtained above and the degenerate two soliton solution reported in Ref. [21, 20] are entirely different as we will see below. To understand the solution (2a)-(2d) further, we analyse ‘$|S^{(l)}|^{2}$’ and ‘$L$’ in the limit $|t|\rightarrow\infty$.
Let us first consider the limit $t\rightarrow-\infty$. In this limit, from the solution (2a)-(2d) under the choice $k_{1I}>l_{1I}$ with $\eta_{1R}\approx 0$ and $\xi_{1R}\rightarrow-\infty$, the expressions for ‘$|S^{(l)}|^{2}$’ and ‘$L$’ reduce to
@@
which corresponds to the soliton of the scalar LSRI system (1), involving a single SW component $S^{(1)}$ and a LW component with amplitude $\frac{1}{2}e^{\rho_{1}-\frac{R_{1}}{2}}$ (for $S^{(1)}$) and $2k_{1R}^{2}$ (for $L$), velocity $2k_{1I}$, and central position $\frac{R_{1}}{2k_{1R}}=\frac{1}{2k_{1R}}\ln\frac{-i\sigma_{1}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}+k_{1}^{*})^{2}(k_{1}-k_{1}^{*})}$. Note that the soliton appearing in the LW component (see Eq. (4b)) arises essentially due to the nonlinearity of the system (1), specifically, only the soliton associated with the $S^{(1)}$-component contributes to the formation of soliton in the LW component since $|S^{(2)}|^{2}$ vanishes as specified by Eq. (4a). Then, by considering the same limit $t\rightarrow-\infty$, we deduce the expressions for ‘$|S^{(l)}|^{2}$’ and ‘$L$’ when $\xi_{1R}\approx 0$ and $\eta_{1R}\rightarrow+\infty$. The resultant expressions are given by
@@
If we now take the limit $t\rightarrow\infty$ with $\eta_{1R}\approx 0$ and $\xi_{1R}\rightarrow\infty$, the expression for ‘$|S^{(l)}|^{2}$’ and ‘$L$’ are turned out to be
@@
(i) Positive nonlinearity ($\sigma_{l}>0$): In this case, the regular solution arises when $k_{1I}<0$ and $l_{1I}<0$. Under this condition, the constants, $e^{R_{n}}$, $n=1,2,3$, are strictly positive. This constraint ensures that solitons in the LSRI model (1), with positive nonlinearity, to propagate in the $-x$-direction, as illustrated in Fig. 3 of our earlier paper Ref. [20].
(ii) Negative nonlinearity ($\sigma_{l}<0$): For the choice $\sigma_{l}<0$, obtaining a regular solution requires $e^{R_{n}}$ ($n=1,2,3$), to be strictly positive, which in turn necessitates $k_{1I},l_{1I}>0$. As a result, the solitons propagate in the opposite direction compared to the previous case. That is, solitons in all the components travel to the $+x$-direction. For brevity, we do not illustrate this scenario here.
(iii) Mixed nonlinearity ($\sigma_{1}>0$ $\sigma_{2}<0$ or $\sigma_{1}<0$ $\sigma_{2}>0$): When considering a mixed choice of nonlinearity strengths to obtain a non-singular solution, the constants $e^{R_{n}}$ ($n=1,2,3$), which appear in the denominators of Eqs. (2a)-(2c), must be positive definite. For instance, if $\sigma_{1}>0$ and $\sigma_{2}<0$, a regular solution requires $k_{1I}<0$ and $l_{1I}>0$. On the other hand, if $\sigma_{1}<0$ and $\sigma_{2}>0$, the condition is reversed, that is, $k_{1I}>0$ and $l_{1I}<0$. This implies that the solitons in the two SW components propagate in opposite directions as shown in Fig. 1. However, in the LW component, these solitons appear and exhibit a two-soliton-like oblique collision. This collision scenario closely resembles a resonant interaction between two Kadomtsev-Petviashvili (KP) line-solitons with a short stem structure [52, 53, 54]. A summary of the propagation properties of the ($1,1,2$)-fundamental non-degenerate soliton solution (8a)-(8c) is provided in Table 1.
@@
$S^{(1)}$ component: Amplitude ($A_{1}$) & Velocity ($v_{1}$) $A_{1}=4k_{1R}\sqrt{|k_{1I}|}$; $v_{1}=-2|k_{1I}|$ – soliton moving in $-x$-direction $A_{1}=4k_{1R}\sqrt{k_{1I}}$; $v_{1}=2k_{1I}$ – soliton moving in $+x$-direction $A_{1}$ $v_{1}$: $4k_{1R}\sqrt{|k_{1I}|}$ $-2|k_{1I}|$ (or $4k_{1R}\sqrt{k_{1I}}$ $2k_{1I}$) – soliton moving in $-x$ (or $+x$)-direction
$S^{(2)}$ component: Amplitude ($A_{2}$) & Velocity ($v_{2}$) $A_{2}=4l_{1R}\sqrt{|l_{1I}|}$; $v_{2}=-2|l_{1I}|$ – soliton moving in $-x$-direction $A_{2}=4l_{1R}\sqrt{l_{1I}}$; $v_{2}=2l_{1I}$ soliton moving in $+x$-direction $A_{2}$ $v_{2}$: $4l_{1R}\sqrt{|l_{1I}|}$ $-2|l_{1I}|$ (or $4l_{1R}\sqrt{l_{1I}}$ $2l_{1I}$) – soliton moving in $-x$ (or $+x$)-direction
LW component ($L$) Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like overtaking-collision in the $-x$-direction. Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like overtaking-collision in the $+x$-direction. Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like head-on collision.
@@ 2.2 $(1,1,1)$ Nondegenerate one-soliton solution
SW components $S^{(1)}$ $S^{(2)}$: Amplitudes ($A_{1}$ $A_{2}$) & Velocity ($v_{1}$) $S^{(1)}\rightarrow$ $A_{1}=4k_{1R}\sqrt{|k_{1I}|}$; $S^{(2)}\rightarrow$ $A_{2}=4l_{1R}\sqrt{|k_{1I}|}$; $v_{1}=-2|k_{1I}|$ – soliton moving in $-x$-direction $S^{(1)}\rightarrow$ $A_{1}=4k_{1R}\sqrt{k_{1I}}$; $S^{(2)}\rightarrow$ $A_{2}=4l_{1R}\sqrt{k_{1I}}$; $v_{1}=2k_{1I}$ – soliton moving in $+x$-direction singular soliton
@@
(i) Positive nonlinearity ($\sigma_{l}>0$): In this case, the regular solution arises when $k_{1I}<0$. Under this condition, the constants, $e^{R_{n}}$, $n=1,2,3$, should be strictly positive. This requirement guarantees that solitons in the LSRI model (1) with positive nonlinearity propagates in the $-x$-direction [20].
(ii) Negative nonlinearity: For the choice $\sigma_{l}<0$, to obtain a regular solution, the quantity $e^{R_{n}}$ ($n=1,2,3$), should be strictly positive, which requires $k_{1I}>0$. Therefore, the solitons propagate to the $+x$-direction.
(iii) Mixed nonlinearity: When a mixed choice of nonlinearity is considered, that is $\sigma_{1}>0$ $\sigma_{2}<0$ or $\sigma_{1}<0$ $\sigma_{2}>0$, the ($1,1,1$) non-degenerate soliton solution (9a)-(9c) always admits singular soliton. We have summarized the propagation properties of the ($1,1,1$)-fundamental non-degenerate soliton solution (9a)-(9c) in Table 2.

### v2

@@ 1 Introduction
In addition to the above, rogue waves, a large localized wave structure that appears and vanishes without a trace [47], and breathers, which are localized breathing structures, have also been investigated in the LSRI system both in the $(1+1)$- and $(2+1)$-dimensional settings, from single-component to multicomponent systems, revealing interesting dynamical patterns [31, 32, 33, 34, 35, 36, 37, 38, 39]. It should be noted that the variants of the LSRI model, such as the Newell (N) type LSRI system [40] and the generalized or N-YO type LSRI system [41], have been proposed. The various nonlinear wave solutions for these models have been reported in the literature [42, 43, 44, 45, 46]. Considering all these developments, in this paper, we intend to investigate the recently proposed non-degenerate vector solitons [48, 49, 50, 51, 52, 53], characterized by non-identical wave numbers, and their several novel features in the following two-component LSRI system
@@ 2 Non-degenerate fundamental soliton solution and its special features
$\hskip-14.22636pte^{R_{1}}=-\frac{i\sigma_{1}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}+k_{1}^{*})^{2}(k_{1}-k_{1}^{*})},~e^{R_{2}}=-\frac{i\sigma_{2}|\alpha_{1}^{(2)}|^{2}}{2(l_{1}+l_{1}^{*})^{2}(l_{1}-l_{1}^{*})},$
$\hskip-14.22636pte^{\Delta_{1}^{(1)}}=-\frac{i\sigma_{2}\alpha_{1}^{(1)}|\alpha_{1}^{(2)}|^{2}(k_{1}-l_{1})}{2(k_{1}+l_{1}^{*})(l_{1}-l_{1}^{*})(l_{1}+l_{1}^{*})^{2}},~e^{\Delta_{1}^{(2)}}=\frac{i\sigma_{1}\alpha_{1}^{(2)}|\alpha_{1}^{(1)}|^{2}(k_{1}-l_{1})}{2(k_{1}^{*}+l_{1})(k_{1}-k_{1}^{*})(k_{1}+k_{1}^{*})^{2}},$
$\hskip-14.22636pte^{R_{3}}=-\frac{\sigma_{1}\sigma_{2}|\alpha_{1}^{(1)}|^{2}|\alpha_{1}^{(2)}|^{2}|k_{1}-l_{1}|^{2}}{4|k_{1}+l_{1}^{*}|^{2}(k_{1}-k_{1}^{*})(l_{1}-l_{1}^{*})(k_{1}+k_{1}^{*})^{2}(l_{1}+l_{1}^{*})^{2}}=e^{R_{1}+R_{2}+\Lambda},~e^{\Lambda}=\frac{|k_{1}-l_{1}|^{2}}{|k_{1}+l_{1}^{*}|^{2}}.$ (2d)
@@
$\hskip-2.84544ptS^{(l)}(x,t)=\frac{\alpha_{1}^{(l)}e^{\eta_{1}}}{1+e^{\eta_{1}+\eta_{1}^{*}+R}}=2A_{l}k_{1R}\sqrt{k_{1I}}e^{i(\eta_{1I}+\frac{\pi}{2})}\mbox{ sech}(\eta_{1R}+\frac{R}{2}),~A_{l}=\alpha_{1}^{(l)}(\sum_{l=1}^{2}\sigma_{l}|\alpha_{1}^{(l)}|^{2})^{-\frac{1}{2}},$ (3a)
$\hskip-2.84544ptL(x,t)=2\frac{\partial^{2}}{\partial x^{2}}\ln(1+e^{\eta_{1}+\eta_{1}^{*}+R})=2k_{1R}^{2}\mbox{ sech}^{2}(\eta_{1R}+\frac{R}{2}),~R=\ln\frac{-\sum_{l=1}^{2}\sigma_{l}|\alpha_{1}^{(l)}|^{2}}{16k_{1R}^{2}k_{1I}},~l=1,2.$ (3b)
The above conventional vector bright soliton is described by only three arbitrary complex parameters, $\alpha_{1}^{(l)}$, $l=1,2$, and $k_{1}$. We remark that one can also obtain the solution (2a)-(2d) from the solution of degenerate two-solitons of the kind (3a)-(3b) in a non-standard way. The degenerate two-soliton solution can be deduced from Eqs. (25a)-(25d)) by setting $N=2$ and by restricting $l_{1}=k_{1}$ ($\xi_{1}=\eta_{1}$) and $l_{2}=k_{2}$ ($\xi_{2}=\eta_{2}$). By substituting $\alpha_{2}^{(1)}=\alpha_{1}^{(2)}=0$ in such degenerate two-soliton solution and renaming the constants, $\alpha_{2}^{(2)}$ and $k_{2}$ as $\alpha_{1}^{(2)}$ and as $l_{1}$, respectively, the solution (2a)-(2d) can be deduced. We have discussed this point in our earlier works in the case of Manakov system [49] and in the case of LSRI system (1) with $\sigma_{1}=\sigma_{2}=1$ [20] . However, in principle, it is possible to derive the non-degenerate soliton solution (2a)-(2d) systematically, see Appendix A of the present work and the methodology described in our earlier work [48, 49, 50, 51, 53, 52]. The dynamical features of such a new class of vector soliton solution (2a)-(2d) obtained above and the degenerate two soliton solution reported in Ref. [21, 20] are entirely different as we will see below. To understand the solution (2a)-(2d) further, we analyse $|S^{(l)}|^{2}$’ and ‘$L$’ in the limit $|t|\rightarrow\infty$.
Let us first consider the limit $t\rightarrow$-$\infty$. In this limit, from the solution (2a)-(2d) under the choice $k_{1I}>l_{1I}$ with $\eta_{1R}\approx 0$ and $\xi_{1R}\rightarrow$-$\infty$, the expressions for $|S^{(l)}|^{2}$’ and ‘$L$’ reduce to
@@
which corresponds to the soliton of the scalar LSRI system (1), involving a single SW component $S^{(1)}$ and a LW component with amplitude $\frac{1}{2}e^{\rho_{1}-\frac{R_{1}}{2}}$ (for $S^{(1)}$) and $2k_{1R}^{2}$ (for $L$), velocity $2k_{1I}$, and central position $\frac{R_{1}}{2k_{1R}}=\frac{1}{2k_{1R}}\ln\frac{-i\sigma_{1}|\alpha_{1}^{(1)}|^{2}}{2(k_{1}+k_{1}^{*})^{2}(k_{1}-k_{1}^{*})}$. Note that the soliton appearing in the LW component (see Eq. (4b)) arises essentially due to the nonlinearity of the system (1), specifically, only the soliton associated with the $S^{(1)}$-component contributes to the formation of soliton in the LW component since $|S^{(2)}|^{2}$ vanishes as specified by Eq. (4a). Then, by considering the same limit $t\rightarrow$-$\infty$, we deduce the expressions for $|S^{(l)}|^{2}$’ and ‘$L$’ when $\xi_{1R}\approx 0$ and $\eta_{1R}\rightarrow+\infty$. The resultant expressions are given by
@@
If we now take the limit $t\rightarrow\infty$ with $\eta_{1R}\approx 0$ and $\xi_{1R}\rightarrow\infty$, the expression for $|S^{(l)}|^{2}$’ and ‘$L$’ are turned out to be
@@
(i) Positive nonlinearity ($\sigma_{l}>0$): In this case, the regular solution arises when $k_{1I}<0$ and $l_{1I}<0$. Under this condition, the constants, $e^{R_{n}}$, $n=1,2,3$, are strictly positive. This constraint ensures that solitons in the LSRI model (1), with positive nonlinearity, to propagate in the x-x-direction, as illustrated in Fig. 3 of our earlier paper Ref. [20].
(ii) Negative nonlinearity ($\sigma_{l}<0$): For the choice $\sigma_{l}<0$, obtaining a regular solution requires $e^{R_{n}}$ ($n=1,2,3$), to be strictly positive, which in turn necessitates $k_{1I},l_{1I}>0$. As a result, the solitons propagate in the opposite direction compared to the previous case. That is, solitons in all the components travel to the +x+x-direction. For brevity, we do not illustrate this scenario here.
(iii) Mixed nonlinearity ($\sigma_{1}>0$ & $\sigma_{2}<0$ or $\sigma_{1}<0$ & $\sigma_{2}>0$): When considering a mixed choice of nonlinearity strengths to obtain a non-singular solution, the constants $e^{R_{n}}$ ($n=1,2,3$), which appear in the denominators of Eqs. (2a)-(2c), must be positive definite. For instance, if $\sigma_{1}>0$ and $\sigma_{2}<0$, a regular solution requires $k_{1I}<0$ and $l_{1I}>0$. On the other hand, if $\sigma_{1}<0$ and $\sigma_{2}>0$, the condition is reversed, that is, $k_{1I}>0$ and $l_{1I}<0$. This implies that the solitons in the two SW components propagate in opposite directions as shown in Fig. 1. However, in the LW component, these solitons appear and exhibit a two-soliton-like oblique collision. This collision scenario closely resembles a resonant interaction between two Kadomtsev-Petviashvili (KP) line-solitons with a short stem structure [54, 55, 56]. A summary of the propagation properties of the ($1,1,2$)-fundamental non-degenerate soliton solution (8a)-(8c) is provided in Table 1.
@@
$S^{(1)}$ component: Amplitude ($A_{1}$) & Velocity ($v_{1}$) $A_{1}=4k_{1R}\sqrt{|k_{1I}|}$; $v_{1}=-2|k_{1I}|$ – soliton moving in x-x-direction $A_{1}=4k_{1R}\sqrt{k_{1I}}$; $v_{1}=2k_{1I}$ – soliton moving in +x+x-direction $A_{1}$ & $v_{1}$: $4k_{1R}\sqrt{|k_{1I}|}$ & $2|k_{1I}|$ (or $4k_{1R}\sqrt{k_{1I}}$ & $2k_{1I}$) – soliton moving in x-x (or +x+x)-direction
$S^{(2)}$ component: Amplitude ($A_{2}$) & Velocity ($v_{2}$) $A_{2}=4l_{1R}\sqrt{|l_{1I}|}$; $v_{2}=-2|l_{1I}|$ – soliton moving in x-x-direction $A_{2}=4l_{1R}\sqrt{l_{1I}}$; $v_{2}=2l_{1I}$ soliton moving in +x+x-direction $A_{2}$ & $v_{2}$: $4l_{1R}\sqrt{|l_{1I}|}$ & $2|l_{1I}|$ (or $4l_{1R}\sqrt{l_{1I}}$ & $2l_{1I}$) – soliton moving in x-x (or +x+x)-direction
LW component ($L$) Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like overtaking-collision in the x-x-direction. Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like overtaking-collision in the +x+x-direction. Amplitude and velocity: R.P and I.P of $k_{1}$ and $l_{1}$, respectively. 2-soliton like head-on collision.
@@ 2.2 $(1,1,1)$ Nondegenerate one-soliton solution
SW components $S^{(1)}$ & $S^{(2)}$: Amplitudes ($A_{1}$ & $A_{2}$) & Velocity ($v_{1}$) $S^{(1)}\rightarrow$ $A_{1}=4k_{1R}\sqrt{|k_{1I}|}$; $S^{(2)}\rightarrow$ $A_{2}=4l_{1R}\sqrt{|k_{1I}|}$; $v_{1}=-2|k_{1I}|$ – soliton moving in x-x-direction $S^{(1)}\rightarrow$ $A_{1}=4k_{1R}\sqrt{k_{1I}}$; $S^{(2)}\rightarrow$ $A_{2}=4l_{1R}\sqrt{k_{1I}}$; $v_{1}=2k_{1I}$ – soliton moving in +x+x-direction singular soliton
@@
(i) Positive nonlinearity ($\sigma_{l}>0$): In this case, the regular solution arises when $k_{1I}<0$. Under this condition, the constants, $e^{R_{n}}$, $n=1,2,3$, should be strictly positive. This requirement guarantees that solitons in the LSRI model (1) with positive nonlinearity propagates in the x-x-direction [20].
(ii) Negative nonlinearity: For the choice $\sigma_{l}<0$, to obtain a regular solution, the quantity $e^{R_{n}}$ ($n=1,2,3$), should be strictly positive, which requires $k_{1I}>0$. Therefore, the solitons propagate to the +x+x-direction.
(iii) Mixed nonlinearity: When a mixed choice of nonlinearity is considered, that is $\sigma_{1}>0$ & $\sigma_{2}<0$ or $\sigma_{1}<0$ & $\sigma_{2}>0$, the ($1,1,1$) non-degenerate soliton solution (9a)-(9c) always admits singular soliton. We have summarized the propagation properties of the ($1,1,1$)-fundamental non-degenerate soliton solution (9a)-(9c) in Table 2.
We have also verified the stability of the obtained $(1,1,1)$-non-degenerate one soliton solution (9a)-(9c) by adopting the split-step Fourier method numerically by introducing random noise as a perturbation in the interval [1,1][-1,1]. For this analysis, we have considered the double-hump profiles in all three components as the initial condition with spatial and temporal step sizes $\Delta x=0.1$ and $\Delta t=0.001$, respectively. The computational domain is chosen as [30,30][-30,30] for $x$ and $[0,5]$ for $t$. The numerical results are presented in Fig. 3, which show that the double-hump profile of the non-degenerate soliton persists even under a $5\%$ of random perturbation. This confirms that the obtained non-degenerate solitons are stable against perturbations.
Figure 3: Panels (a1)-(a3) illustrate the propagation of the double-hump profile of the non-degenerate soliton in the absence of perturbation, while panels (b1)-(b3) demonstrate the stability of the same double-hump soliton under 5% random noise perturbation.